### v1

# Spatially Regularized Super-Resolved Constrained Spherical Deconvolution (SR2-CSD) of Diffusion MRI Data

Abstract
Constrained Spherical Deconvolution (CSD) is crucial for estimating white matter fiber orientations using diffusion MRI data. A relevant parameter in CSD is the maximum order $l_{max}$ used in the spherical harmonics series, influencing the angular resolution of the Fiber Orientation Distributions (FODs). Lower $l_{max}$ values produce smoother and more stable estimates, but result in reduced angular resolution. Conversely, higher $l_{max}$ values, as employed in the Super-Resolved CSD variant, are essential for resolving narrow inter-fiber angles but lead to spurious lobes due to increased noise sensitivity. To address this issue, we propose a novel Spatially Regularized Super-Resolved CSD (SR2-CSD) approach, incorporating spatial priors into the CSD framework. This method leverages spatial information among adjacent voxels, enhancing the stability and noise robustness of FOD estimations. SR2-CSD facilitates the practical use of Super-Resolved CSD by including a J-invariant auto-calibrated total variation FOD denoiser. We evaluated the performance of SR2-CSD against standard CSD and Super-Resolved CSD using phantom numerical data and various real brain datasets, including a test-retest sample of six subjects scanned twice. In phantom data, SR2-CSD outperformed both CSD and Super-Resolved CSD, reducing the angular error (AE) by approximately half and the peak number error (PNE) by a factor of three across all noise levels considered. In real data, SR2-CSD produced more continuous FOD estimates with higher spatial-angular coherency. In the test-retest sample, SR2-CSD consistently yielded more reproducible estimates, with reduced AE, PNE, mean squared error, and increased angular correlation coefficient between the FODs estimated from the two scans for each subject.
Keywords diffusion MRI  $\cdot$ spherical deconvolution  $\cdot$ spatial regularization  $\cdot$ white matter  $\cdot$ quadratic programming
1 Introduction
Diffusion Magnetic Resonance Imaging (dMRI) is sensitive to the microscopic displacements of water molecules in biological tissues. This capability makes it helpful for non-invasively exploring the complex tissue architecture of the brain’s white matter (WM). Numerous techniques have been developed to estimate local WM fiber orientations [1], including diffusion tensor imaging [2], diffusion spectrum imaging [3, 4], q-ball imaging [5, 6, 7, 8, 9], diffusion orientation transform [10, 11], higher-order tensor models [12, 13, 14], and multi-tensor fitting algorithms [15]. Each technique offers distinct advantages and limitations. However, spherical deconvolution (SD) methods [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] have emerged as the leading approach for estimating voxel-wise fiber orientation distribution (FOD). These methods provide superior resolution for detecting fiber crossings with small inter-fiber angles, utilizing single-shell [20] or multi-shell [36] high angular resolution diffusion imaging (HARDI) data [15].
Among the various types of parametric and non-parametric SD methods, constrained spherical deconvolution (CSD) [20], as implemented in MRtrix3 [37] and Diffusion Imaging in Python (DIPY) [38] software packages, is more commonly used in clinical applications and research. This state-of-the-art method employs orthogonal spherical harmonics (SH) functions to represent the FOD, providing a compact, efficient and continuous representation of functions defined on the surface of a sphere. Notably, CSD uses a non-negativity constraint on the FOD to mitigate the presence of nonphysical negative FOD lobes. This constraint provides supplementary information facilitating the implementation of super-resolution: a CSD variant known as Super-CSD, wherein the number of estimated spherical harmonics (SH) coefficients exceeds the number of dMRI measurements [20].
Despite the significant advancements introduced by CSD [20], several limitations remain. A crucial parameter affecting the quality of the estimated FOD is the maximum order $l_{max}$ used in the SH series expansion, which determines the maximum achievable angular resolution. A low value (e.g., $l_{max}<8$) can produce smooth and stable estimates from noisy dMRI data by discarding higher frequency components. Conversely, using a higher $l_{max}$ in Super-CSD can yield sharper FODs with increased angular resolution, essential for resolving narrower inter-fiber angles. Although initial results with phantom dMRI data [39] for Super-CSD with $l_{max}=16$ were promising, CSD with $l_{max}=8$ remains the recommended method for real brain data [20]. This suggests that the non-negativity constraint alone may not enable robust angular super-resolution. The deconvolution inverse problem is ill-conditioned and susceptible to noise contamination, meaning the increased angular resolution provided by Super-CSD in some regions does not compensate for its inferior estimation stability. This often results in FODs with spurious lobes and less coherent orientations across adjacent voxels, as noted in [20].
To address this limitation, we introduce a new Spatially Regularized Super-Resolved CSD (SR2-CSD) approach that leverages the redundancy and spatial correlations of FODs among adjacent voxels as additional prior information, alongside the non-negativity constraint. We hypothesize that incorporating spatial information will enhance the proposed method’s stability and noise robustness, enabling the benefits of angular super-resolution while minimizing the appearance of spurious lobes. Consequently, we anticipate that SR2-CSD will produce more spatially coherent and reproducible FOD estimates than the standard CSD.
2 Theory
[…]
$\boldsymbol{u}=\boldsymbol{P}\boldsymbol{f}_{i}.$ (1)
Then, a regularization matrix $\boldsymbol{L}_{i}$ is constructed as follows:
$L_{m,n}=\begin{cases}P_{m,n},&\text{if $u_{m}<\tau$ }\\ 0,&\text{if $u_{m}\geq\tau$,}\end{cases}$ (2)
[…]
$\boldsymbol{f}_{i+1}=argmin\{||\boldsymbol{Af}_{i}-\boldsymbol{s}||^{2}+\lambda||\boldsymbol{L}_{i}\boldsymbol{f}_{i}||^{2}\},$ (3)
[…]
$\boldsymbol{f}_{i+1}=(\boldsymbol{A}^{T}\boldsymbol{A}+\lambda\boldsymbol{L}_{i}^{T}\boldsymbol{L}_{i})^{-1}\boldsymbol{A}^{T}\boldsymbol{s}.$ (4)
The order $l_{max}$ needed to build the matrix $\boldsymbol{A}$ is typically chosen such that the number of SH coefficients $p=(l_{max}+1)(l_{max}+2)/2$ to be estimated is less than or equal to the number of measurements $n$. However, due to the prior information introduced by the non-negativity constraint, it is possible to estimate a number of SH coefficients $p>n$, allowing the FOD to be represented using higher-order SH coefficients. This variant is known as Super-CSD [20]. Despite the superior performance of Super-CSD in resolving narrower fiber crossings, this variation is more sensitive to noise and produces less stable estimations. Therefore, $l_{max}=8$ is employed in practice to balance the trade-off between CSD angular resolution and noise robustness.
We noted that although the original SSST-CSD method was implemented using this approach [20], the implementation in MRtrix3 [37] was slightly modified to include a new regularization term for improving the estimation stability:
$\boldsymbol{f}_{i+1}=(\boldsymbol{A}^{T}\boldsymbol{A}+\lambda_{1}\boldsymbol{L}_{i}^{T}\boldsymbol{L}_{i}+\lambda_{2}\boldsymbol{I})^{-1}\boldsymbol{A}^{T}\boldsymbol{s}.$ (5)
On the other hand, in the multi-shell, multi-tissue CSD (MSMT-CSD) version [36], the cost function in Eq. (3) is replaced by
$\begin{split}\boldsymbol{\tilde{f}}=argmin\{||\boldsymbol{Af}-\boldsymbol{s}||^{2}\},\\ \text{subject to }\boldsymbol{Pf}\geq\boldsymbol{0},\end{split}$ (6)
which can be written as a Quadratic Programming (QP) problem, strictly enforcing the non-negativity constraint on the FOD amplitudes:
$\begin{split}\boldsymbol{\tilde{f}}=argmin\{\frac{1}{2}\boldsymbol{f}^{T}\boldsymbol{Q}\boldsymbol{f}+\boldsymbol{c}^{T}\boldsymbol{f}\}\\ \text{subject to }\boldsymbol{Kf}\leq\boldsymbol{q},\end{split}$ (7)
where $\boldsymbol{K}$=$-\boldsymbol{P}$, $\boldsymbol{q}=\boldsymbol{0}$, $\boldsymbol{Q}$=$\boldsymbol{A}^{T}\boldsymbol{A}$, and $\boldsymbol{c}^{T}=-\boldsymbol{s}^{T}\boldsymbol{A}$.
In this study, the CSD and Super-CSD estimates corresponding to the SSST-CSD method were computed using Eq. (5), employing the default parameters in MRtrix3 [37]: $\lambda_{1}=k_{1}*c$, $\lambda_{2}=k_{2}*d$, where $c=max(\boldsymbol{A}^{T}\boldsymbol{A})$, $d=2*10^{-4}*c$, and $k_{1}$, $k_{2}$ and $\tau$ are the input parameters with default values of $k_{1}=k_{2}=1$ and $\tau=0$. Conversely, the estimates corresponding to MSMT-CSD were computed using Eq. (7) utilizing the quadprog [40] solver available at https://github.com/qpsolvers/qpsolvers.
2.1 Spatially Regularized Super-Resolved CSD (SR2-CSD)
Nerve fibers should exhibit spatial continuity across WM voxels, as they do not abruptly begin and end their trajectories within the WM. Therefore, utilizing information about the estimates in neighboring voxels can lead to more accurate results. The standard CSD formulation does not incorporate spatial information, as it performs the estimation independently for each voxel. This study introduces a novel method, SR2-CSD, which incorporates a spatial prior along with the non-negativity constraint. As demonstrated in our experiments, this additional information helps stabilize the Super-CSD technique, enabling the practical use of larger $l_{max}$ values. This results in estimates with higher angular resolution and a reduction in the number of spurious FOD lobes compared to standard CSD.
The SR2-CSD inverse problem incorporates spatial information between adjacent voxel solutions as prior information to enhance denoising and promote spatial continuity of the estimated FODs. It is formulated in the following general form:
$\begin{split}\boldsymbol{\tilde{f}}=argmin\{||\boldsymbol{Af}-\boldsymbol{s}||^{2}+\rho||\boldsymbol{f-f_{0}}||^{2}\},\\ \text{subject to }\boldsymbol{Pf}\geq\boldsymbol{0},\end{split}$ (8)
where $\boldsymbol{f}_{0}$ denotes a prior estimate of the FOD SH coefficients, computed by incorporating spatial information, and the non-negative parameter $\rho$ controls the weighting between the data-fitting and regularization terms. Notably, Eq. (8) can be reformulated as a QP problem similar to Eq. (7), but with parameters $\boldsymbol{K}$=$-\boldsymbol{P}$, $\boldsymbol{q}=\boldsymbol{0}$, $\boldsymbol{Q}$=$\boldsymbol{A}^{T}\boldsymbol{A}+\rho\boldsymbol{I}$, and $\boldsymbol{c}^{T}=-\boldsymbol{s}^{T}\boldsymbol{A}-\rho\boldsymbol{f}_{0}^{T}$.
Various strategies could be implemented to obtain $\boldsymbol{f}_{0}$. In this study, we computed it as a Total Variation (TV) denoised version of the solution produced by the Super-CSD technique. TV denoising is a highly effective spatial filter, as it simultaneously removes noise in flat regions while preserving edges. Importantly, the smoothing is primarily applied adaptively along the dominant directions of the image structures, preserving fine details. This property is particularly advantageous for CSD, as it promotes smoothness and continuity along individual tracts, while minimizing partial volume contamination between adjacent tracts [32]. Note that this formulation can be used in both SSST-CSD [20] or MSMT-CSD [36] variants.
In practice, the TV filter [41] is applied independently to each 3D image formed by the FOD SH coefficients of a given degree and order. The strength of the filtering, defined by $\lambda=K\sigma$, varies depending on the noise standard deviation $\sigma$ affecting each SH 3D map. This noise standard deviation is estimated using a robust wavelet-based estimator [42], implemented in the estimate_sigma function available in the scikit-image Python package [43]. In the results section, we demonstrate the impact of the solution as a function of $K$, which is kept constant for all SH coefficients. However, since the optimal $K$ should depend on the complexity of the underlying anatomy of the image, it is generally not feasible to define a constant $K$ value that is optimal across different datasets and phantoms. To address this, we employ the J-invariance principle [44, 45] (see Section 2.2) to calibrate the parameter $K$ automatically. In our implementation, TV denoising is implemented using the Chambolle method [46], available in the scikit-image package. The smoothed FOD SH coefficients are then projected onto the constrained non-negative set, ensuring all the SH coefficients produce non-negative FOD amplitudes. This projection involves calculating the FOD amplitudes, setting negative values to zero, and recomputing the SH coefficients from the thresholded FODs. The resulting prior $\boldsymbol{f}_{0}$ is then used to solve the QP problem in Eq. (8). The regularization parameter is chosen as $\rho=1$. The pipeline is implemented in the DIPY software [38].
2.2 J-Invariance for Automatic TV Denoiser Calibration
The automatic TV denoiser calibration uses a self-supervised approach based on the Noise2Self algorithm [44]. A denoising function is considered J-invariant if the value of each pixel in the denoised output does not depend on the value of the same pixel in the noisy input image. If a function is not J-invariant, a masking procedure can be applied to make it J-invariant [44]. For a J-invariant denoiser, the self-supervised loss is equivalent to the ground-truth loss, up to a constant. This means that the difference between the denoised image and the noisy image can be used as a proxy for the difference between the denoised image and the original clean image. Therefore, we can calibrate a J-invariant denoiser using self-supervised loss, even without clean reference images. In our experiments, we searched for the optimal TV filtering strength $K$ within the range [0, 5].
3 Materials and Methods
The performance of CSD, Super-CSD, and SR2-CSD is compared quantitatively and visually using synthetic data from a numerical phantom. Additionally, the impact of the proposed technique is evaluated on real brain dMRI data. The following subsections provide more details about the employed phantom, real data, acquisition parameters, evaluation metrics, and data preprocessing.
3.1 HARDI Phantom
The synthetic dMRI phantom from the HARDI Reconstruction Challenge 2013, organized as part of the IEEE International Symposium on Biomedical Imaging 2013, was utilized to evaluate the algorithms. The dataset is designed to simulate HARDI data by employing a dMRI model that accounts for intra- and extra-cellular water. The phantom consists of 27 fiber bundles connecting different regions of a 3D image with dimensions of 50x50x50 voxels. It features various configurations, including branching, crossing, and kissing, along with diverse fiber bundle radii and geometric properties. Additionally, the phantom incorporates isotropic compartments to mimic the CSF partial volume contamination observed near ventricles in real brain images [47]. The dMRI data was generated using a b-value of 3000 s/mm2 and $n$=64 diffusion directions, along with one additional $b=0$ s/mm2 image. The dataset can be downloaded from the official webpage: http://hardi.epfl.ch/static/events/2013_ISBI/.
Rician noise was added to the noise-free dMRI signals $E$ as follows:
$E_{noisy}=\sqrt{(E+\eta_{1})^{2}+\eta_{2}^{2}},$ (9)
where $\eta_{1}$ and $\eta_{2}$ denote noise vectors sampled from zero-mean, independent Gaussian distributions with variance $\sigma^{2}$. The noise was added using three signal-to-noise ratio (SNR) levels: SNR=10, 15, and 30, defined as SNR=$\frac{S_{0}}{\sigma}$, with $S_{0}$ representing the non-diffusion-weighted image for $b=0$ s/mm2. For each SNR level, 10 datasets were generated using different noise realizations to assess the stability of the methods.
3.2 Sherbrooke Data
A multi-shell HARDI dataset was obtained from a healthy human volunteer by the Sherbrooke Connectivity Imaging Lab (SCIL). This dataset includes dMRI images acquired at three b-values (1000, 2000, and 3500 s/mm2), each with 64 gradient directions. Additionally, one non-diffusion-weighted image was collected. The spatial resolution is 2x2x2 mm3. The dataset is available as part of the DIPY software at https://docs.dipy.org/stable/user_guide/data.html.
For analysis, the single-shell single-tissue CSD method uses the subset of data with $b$=3500 s/mm2. All three b-values are utilized to estimate the multi-shell multi-tissue CSD.
3.3 Test-Retest dMRI Data
We assessed method reproducibility using data from the Center for Cognitive and Neurobiological Imaging at Stanford University, obtained with a 3T GE Discovery MR750 MRI system [48]. The dataset consists of scans from six healthy male participants, each undergoing two consecutive scans. Images were acquired at a spatial resolution of 1.5x1.5x1.5 mm3. The dataset includes 10 $b0$ images and 96 diffusion-weighting gradient directions with a b-value of 2000 s/mm2.
3.4 Evaluation Metrics
This subsection introduces four voxel-wise evaluation metrics used in our study, with their averaged values reported across the extracted white matter voxels.
  • Mean Squared Error (MSE): The FOD SH coefficients estimated from two scans are normalized by their total "energies" voxel-by-voxel using Eq. (10) to eliminate scaling differences.
    $SH_{norm}[j]=\frac{SH[j]}{\sqrt{\sum_{i=1}^{p}SH[i]^{2}}},$ (10)
    where the index $j=(l,m)$ encodes the degree and order of the SH coefficients. Then, the MSE between the normalized coefficients $SH_{est1}$ and $SH_{est2}$ is calculated as:
    $MSE=\frac{1}{p}\sum_{j=1}^{p}(SH_{est1}[j]-SH_{est2}[j])^{2}.$ (11)
  • The Angular Correlation Coefficient (ACC) is calculated per voxel as follows [49]:
    $ACC=\frac{\sum_{j=1}^{p}SH_{est1}[j]SH_{est2}[j]^{*}}{\sqrt{\sum_{j=1}^{p}|SH_{est1}[j]|^{2}}\sqrt{\sum_{j=1}^{p}|SH_{est2}[j]|^{2}}},$ (12)
    where $SH_{est2}[j]^{*}$ is the complex conjugate of $SH_{est2}[j]$.
  • The Angular Error (AE) between two estimated FODs is evaluated by measuring the angular separation (in degrees) between their respective fiber orientations $\vec{d}_{est1}$ and $\vec{d}_{est2}$, as outlined in [15, 29, 50]:
    $AE=\arccos(\vec{d}_{est1}\cdot\vec{d}_{est2})\cdot\frac{180}{\pi}.$ (13)
    Here, the fiber orientations are unit vectors derived from the FOD peaks. Initially, the fiber orientations from both FODs are paired based on their proximity, such that $\vec{d}_{est2}$ represents the FOD peak from the second FOD that is closest to the FOD peak $\vec{d}_{est1}$ extracted from the first FOD. The resultant value reported for each voxel is the mean AE computed across all fiber orientations, as described in [29].
  • The Peak Number Error (PNE) quantifies the normalized disparity between the estimated numbers of paired peaks in the FODs, as defined in [29]:
    $PNE=\frac{|M_{est1}-M_{est2}|}{M_{est1}},$ (14)
[…]
For the HARDI dataset, the ground-truth fiber directions are known. Hence, to calculate the AE and PNE metrics, we substitute $\vec{d}_{est1}$ and $M_{est1}$ with their corresponding ground-truth values $\vec{d}_{true}$ and $M_{true}$, respectively.
Conversely, in the real data reproducibility study, all four metrics were computed using the FODs estimated from the two scan sessions. Additionally, since the AE and PNE metrics are not necessarily symmetric, we report their average values by considering the estimates from both the first and second scan sessions as reference values.
3.5 Preprocessing
Before computing the FODs, all dMRI data used in this study, including the HARDI phantom, Sherbrooke dataset, and the Test-Retest dMRI dataset, underwent denoising using the state-of-the-art Marchenko–Pastur Principal Component Analysis (MPPCA) method [51, 52, 53].
The Test-Retest data has already been preprocessed for motion, and measurements of the B0 magnetic field were performed to mitigate the effects of EPI spatial distortions, as detailed in the Stanford Digital Repository available at https://purl.stanford.edu/rt034xr8593. Despite this preprocessing, the two images collected per subject in this dataset could not be optimally spatially registered using a linear registration method due to residual spatial distortions affecting each image differently. Advanced tools like topup [54, 55] and eddy [56] in FSL could not be applied to further attenuate these distortions because no $b=0$ images were acquired with opposing polarities of the phase-encode blips. Instead, the estimated FODs for each subject from the two scans were spatially normalized using a non-linear registration algorithm, as described by [57], to achieve better spatial agreement.
Figure 1: Comparison of Angular Error (a) and Peak Number Error (b) for the SR2-CSD method with $l_{max}=12$ as a function of the TV filtering strength parameter $K$ in the HARDI Phantom dataset (SNR = 15). Results from CSD with $l_{max}=8$ (blue horizontal line) and Super-CSD with $l_{max}=12$ (orange horizontal line) are shown for reference. The vertical red line indicates the $K$ value determined by the J-invariance calibrated TV denoiser, aligning with the global minima.
The response functions for SSST-CSD and MSMT-CSD experiments were estimated using DIPY. For SSST-CSD, the single-fiber response function was estimated using the response_from_mask_ssst function from the top 200 voxels with the highest fractional anisotropy. For MSMT-CSD, the response functions for WM, GM, and CSF tissue types were estimated using the mask_for_response_msmt function with default parameters as described in the DIPY documentation https://docs.dipy.org/stable/examples_built/reconstruction/reconst_mcsd.html (i.e., wm_fa_thr=0.7, gm_fa_thr=0.3, csf_fa_thr=0.15, gm_md_thr=0.001, csf_md_thr=0.0032).
Table 1: Results from HARDI Phantom for SNR = 10, 15, and 30. Angular Error (AE) and Peak Number Error (PNE) metrics were computed for CSD with $l_{max}=8$, and Super-CSD and SR2-CSD with $l_{max}=12$. The mean and standard deviation were calculated from estimates based on 10 different noise realizations. The optimal results for each SNR are highlighted in bold.
HARDI Data (SNR = 10)
Denoised Data AE (degree) PNE
CSD 21.635 $\pm$ 0.068 0.626 $\pm$ 0.002
Super-CSD 22.225 $\pm$ 0.072 0.656 $\pm$ 0.002
SR2-CSD 10.363 $\pm$ 0.497 0.210 $\pm$ 0.015
HARDI Data (SNR = 15)
Denoised Data AE (degree) PNE
CSD 17.592 $\pm$ 0.073 0.553 $\pm$ 0.002
Super-CSD 17.822 $\pm$ 0.084 0.570 $\pm$ 0.002
SR2-CSD 7.831 $\pm$ 0.133 0.145 $\pm$ 0.004
HARDI Data (SNR = 30)
Denoised Data AE (degree) PNE
CSD 14.531 $\pm$ 0.097 0.468 $\pm$ 0.004
Super-CSD 14.877 $\pm$ 0.068 0.484 $\pm$ 0.003
SR2-CSD 7.624 $\pm$ 0.057 0.140 $\pm$ 0.001
Figure 2: Fiber Orientation Distributions (FODs) estimated in two fiber crossing regions of the HARDI Phantom with SNR = 15 (yellow and orange boxes). The first row shows the ground truth peak orientations, and the second and third row displays the ground truth peaks, and the FODs estimated using CSD with $l_{max}=8$, Super-CSD with $l_{max}=12$, and SR2-CSD with $l_{max}=12$. Three regions are highlighted where SR2-CSD improved the angular resolution (red boxes) and the angular coherency (blue box).
4 Results
4.1 Phantom
In the HARDI dataset, we compared the AE and PNE metrics for CSD with $l_{max}=8$, Super-CSD with $l_{max}=12$, and SR2-CSD with $l_{max}=12$ using the SSST-CSD variant. For SR2-CSD, the prior $\boldsymbol{f_{0}}$ was obtained for different values of the TV filtering strength $K$.
Figure 1 illustrates how the AE and PNE metrics derived from the SR2-CSD method depend on $K$ for data with SNR = 15. Notably, the optimal $K$ found by the self-calibrated TV denoiser (indicated by the vertical red line) produced the best results. Moreover, within the explored range of $K$ values, any solution with $K>0$ produced superior metrics compared to those obtained from CSD and Super-CSD, shown as blue and orange horizontal lines, respectively, for reference.
Table 1 reports the AE and PNE metrics for the HARDI dataset with SNR = 10, 15, and 30. The mean and standard deviation of each metric are reported using ten different noise realizations. The estimates for SR2-CSD correspond to the optimal value of $K$ determined automatically. The SR2-CSD method outperformed both CSD and Super-CSD, approximately reducing the AE by half and the PNE by a factor of three in all cases.
To complement the numerical experiments, Figure 2 depicts the estimated FODs in various regions of interest. CSD with $l_{max}=8$ and $l_{max}=12$ could not resolve some fiber crossings at small inter-fiber angles, whereas SR2-CSD successfully detected these crossings (see the highlighted red boxes in both figures). Furthermore, SR2-CSD enhances spatial angular coherency, producing aligned peak orientations and FOD lobes with intensities that agree more with the ground truth values (see the highlighted blue boxes in Figure 2).
Figure 3: Fiber Orientation Distributions (FODs) estimated from the Sherbrooke dataset using $b$ = 3500 s/mm2 and $n$ = 64 DW directions. The estimated FODs are displayed, from left to right, for CSD with $l_{max}=8$, Super-CSD with $l_{max}=12$, and SR2-CSD with $l_{max}=12$. The red boxes highlight two regions where the SR2-CSD method produced more spatially continuous FODs. The green FOD lobes (corresponding to the superior longitudinal fasciculus) estimated by CSD and Super-CSD are highly attenuated in those regions.
Figure 4: Fiber Orientation Distributions (FODs) estimated from the Sherbrooke dataset in a region of interest including WM, GM, and CSF voxels. The estimated FODs are shown from left to right for CSD with $l_{max}=8$, Super-CSD with $l_{max}=12$, and SR2-CSD with $l_{max}=12$. Results from both the single-shell single-tissue CSD (SSST-CSD, obtained from the data with $b$=3500 s/mm2) and multi-shell multi-tissue CSD (MSMT-CSD, using three b-values: 1000, 2000, 3500 s/mm2) variants are shown.
Table 2: Results from the reproducibility study for CSD with $l_{max}=8$, and Super-CSD and SR2-CSD with $l_{max}=12$. Mean squared error (MSE), angular correlation coefficient (ACC), angular error (AE), and peak number error (PNE) metrics were calculated using the two Fiber Orientation Distribution (FOD) estimates from the two scans per subject. The optimal metrics for each subject are highlighted in bold.
Subject MSE ACC AE PNE Subject MSE ACC AE PNE
Subject 1 Subject 4
CSD 0.265 0.867 19.30 0.188 CSD 0.266 0.871 19.01 0.191
Super-CSD 0.342 0.828 20.68 0.203 Super-CSD 0.337 0.835 20.10 0.208
SR2-CSD 0.180 0.910 17.53 0.184 SR2-CSD 0.173 0.918 16.70 0.184
Subject 2 Subject 5
CSD 0.277 0.861 20.07 0.190 CSD 0.234 0.883 19.75 0.209
Super-CSD 0.362 0.819 21.36 0.197 Super-CSD 0.316 0.842 20.99 0.217
SR2-CSD 0.179 0.912 18.09 0.188 SR2-CSD 0.179 0.910 17.86 0.206
Subject 3 Subject 6
CSD 0.295 0.852 19.65 0.182 CSD 0.325 0.837 20.74 0.194
Super-CSD 0.364 0.817 20.69 0.202 Super-CSD 0.407 0.796 21.74 0.205
SR2-CSD 0.193 0.903 17.33 0.175 SR2-CSD 0.224 0.888 18.83 0.203
Figure 5: Scan-rescan results for Subject 3. Panels (a)-(b): Voxelwise maps of Mean Squared Error (MSE) and Angular Correlation Coefficient (ACC) for CSD with $l_{max}=8$, Super-CSD with $l_{max}=12$, and SR2-CSD with $l_{max}=12$. Panels (c)-(d): Fiber Orientation Distributions (FODs) estimated from the two scans. Two regions of interest are shown, with boxes highlighting areas where SR2-CSD produced more similar FODs.
4.2 Sherbrooke Data
Figure 3 examines a region in an axial slice where various fibers cross, including the first branch of the superior longitudinal fasciculus (SLF I), which connects the superior parietal and superior frontal lobes [58], and the anterior part of the Corpus Callosum (CC). The SLF fibers correspond to the green FOD lobes in the figure, and the blue FOD lobes correspond to the CC. The highlighted red boxes indicate some voxels where SR2-CSD can better detect this crossing, producing more continuous FOD estimates with higher spatial-angular coherency. These results were obtained using the SSST-CSD variant for the data acquired with $b$=3500 s/mm2. The optimal TV filtering strength estimated by the J-invariance approach was $K=0.45$.
FODs from both SSST-CSD and MSMT-CSD versions at the WM-GM-CSF interface in an example region are depicted in Figure 4. For the three methods, results from the SSST-CSD versions appear noisier, as expected, since a single b-value is inefficient in modeling the signal contributions from distinct tissue compartments. In each case, the FODs estimated by SR2-CSD from the WM compartment are smaller outside the WM voxels, where we don’t expect to observe fiber bundles, and show a reduction in spurious lobes.
4.3 Test-Retest dMRI Data
Two FOD reconstructions were obtained for each of the six subjects included in this dataset, from the two scanning sessions. The reproducibility of the evaluated methods is assessed by calculating four similarity metrics between these two reconstructions, as defined in Section 3.4: MSE, ACC, AE, and PNE. Lower MSE, AE, and PNE metrics, and higher ACC metrics indicate more reproducible FOD estimates. Table 2 reports the mean values and standard deviations for the evaluated metrics for each subject, for the CSD with $l_{max}=8$, and Super-CSD and SR2-CSD methods with $l_{max}=12$. All metrics were computed in a WM mask created by selecting all voxels with a fractional anisotropy higher than 0.3. The mean (and standard deviation) of the TV filtering strength found by the J-invariance approach from all subjects was $K$=0.44 (0.08).
Notably, SR2-CSD outperformed the other two methods in all the metrics for all subjects, except for the PNE metric for the sixth subject, where CSD with $l_{max}=8$ obtained the best score, closely followed by SR2-CSD.
Figure 5 shows a coronal view of the MSE and ACC voxel-wise maps for Subject 3, comparing the results from CSD, Super-CSD, and SR2-CSD. Consistent with the results reported in Table 2, SR2-CSD produced more reproducible estimates with the lowest MSE and highest ACC metrics. Figure 5 also displays the FODs estimated from the two scans for Subject 3 in two regions of interest in sagittal and coronal slices. Once again, SR2-CSD produced more similar estimates than CSD and Super-CSD (see the highlighted voxels).
5 Discussion and Conclusion
We propose a novel Spatially Regularized Super-Resolved CSD (SR2-CSD) approach that leverages a spatial prior obtained through Total Variation filtering applied to the fiber orientation distribution estimated by the Super-CSD method. The TV filtering strength is automatically calibrated based on the J-invariance principle, a technique grounded in self-supervised blind denoising.
To evaluate the performance of the SR2-CSD method, we compared it with the standard CSD method using $l_{max}=8$ and Super-CSD with $l_{max}=12$. Our experiments utilized both phantom and various real brain datasets, employing quantitative metrics and qualitative visual assessments for performance analysis. The phantom data estimates were evaluated under different SNR levels and noise realizations, while a test-retest analysis was conducted on real brain data.
In the phantom data, numerical metrics demonstrate that SR2-CSD consistently produced lower angular and peak number errors across three different SNRs (Table 1). Visual comparisons (Figure 2) corroborate these numerical results, showing SR2-CSD’s superior ability to detect fiber crossings with small inter-fiber angles and enhance the spatial-angular coherence of the estimated FODs.
In real data, SR2-CSD generated smoother FOD estimates without sacrificing angular resolution, potentially providing a better representation of continuous fiber tracts. Single-shell single-tissue and multi-shell multi-tissue analyses were compared at the WM-GM-CSF interfaces (Figure 4). Consistent with the findings of [36], single-shell analyses were insufficient to characterize signal contributions from non-WM compartments. Notably, the SR2-CSD’s inherent denoising capability reduced the noisy FODs produced by CSD and Super-CSD in these regions. A similar noise reduction trend was observed in the multi-shell multi-tissue analysis, though all methods produced smoother FODs and presumably fewer spurious lobes.
Reproducibility was assessed using a dataset of six subjects scanned twice. According to four metrics comparing the similarity of the estimated FODs, SR2-CSD produced more reproducible reconstructions. The quantitative analysis in Table 2 was supported by visual inspections of the estimated FODs (Figure 5).
The computation time of the proposed SR2-CSD method is longer than that of Super-CSD, as it requires solving the QP problem twice. The first iteration computes the initial solution used to build the prior, and the second iteration solves the augmented problem using this prior. Notably, the second iteration is faster than the first because the prior helps the solver converge more quickly. Here, we report the computation times (CT) of the different steps for various datasets. For the HARDI data: Super-CSD (first iteration) CT=34 secs, TV denoising CT=32 secs, and SR2-CSD fitting (second iteration) CT=22 secs using an AMD Ryzen 9 5900X-12 Core-Processor. For Sherbrooke data: Super-CSD (first iteration) CT=7 mins 36 secs, TV denoising CT=10 mins 34 secs, and SR2-CSD fitting (second iteration) CT=1 min 28 secs. In practice, the CT of the TV filtering step could be decreased by reducing the range of $K$ values, as the optimal $K$ was in the range from 0.35 to 0.65 for all the real data used in this study.
Before computing the FODs, all dMRI data used in this study—including the HARDI phantom, Sherbrooke data, and Test-Retest dataset—underwent denoising using the MPPCA method [51, 52, 53]. This denoising step was crucial to ensure a fair comparison between the CSD, Super-CSD, and SR2-CSD methods. Without denoising, differences in the performance of these methods could be mainly attributed to the presence of noise and the superior noise reduction of SR2-CSD. We attenuated the noise using a state-of-the-art technique before applying the three methods, enabling a more equitable performance comparison. Notably, when we repeated all experiments using the non-denoised dMRI data, SR2-CSD showed even greater improvements over the other methods compared to the results based on denoised data, as expected (results not shown).
Future studies should explore alternative denoising methods for obtaining the spatial prior needed in SR2-CSD, including machine learning techniques like [59]. Additionally, our two-step analysis—computing the prior and re-estimating the FOD using the prior—could evolve into a more generalized iterative algorithm that minimizes a global multi-voxel cost function. This function would penalize the mean squared error between the measured and predicted data by the deconvolution model and the noise and lack of spatial continuity of the estimated FODs, potentially using a projected gradient descent method. However, we defer this implementation to future work, as incorporating the non-negativity constraint into this variant requires further investigation.
Future research should also evaluate whether the improved FOD estimates from SR2-CSD positively impact the connectivity estimation obtained by tractography methods [60], particularly regarding false positives and false negatives [61], as well as tracking reproducibility and subject-specificity [62]. The proposed method will be made freely available to the community in the DIPY software library [38].
6 Acknowledgments
This work is supported by the Swiss National Science Foundation (SNSF) under grant Nbr 205320__\__204097. Erick J. Canales-Rodríguez was supported by the SNSF Ambizione grant PZ00P2__\__185814.